# Topological Defects in Spiral Wave Chimera States

### v1

@@ Abstract
Chimera states, where coherent and incoherent domains coexist, represent a key self-organization phenomenon in the study of synchronization in complex systems. We introduce a topological analysis method based on winding numbers to characterize the dynamics of spiral wave chimeras in a two-dimensional phase oscillator network. Our investigation reveals distinct scaling laws governing the system’s evolution across the phase lag $\alpha$. Perturbation analysis in the limit $\alpha\to 0$ demonstrates that the incoherent core radius scales linearly with $\alpha$. In contrast, within the stable chimera regime, the average total positive winding number $\overline{n_{+}}$ follows a clear exponential growth law $\overline{n_{+}}=ae^{b\alpha}$. This divergence suggests a physical crossover from geometric core expansion to active topological excitation. Furthermore, we identify a statistical transition in the defect distribution from binomial-like to Poisson-like behavior at a critical threshold $\alpha^{*}$, which we interpret as a shift from a constrained regime to an unconstrained state analogous to a BKT binding-unbinding transition. These results indicate that topological defects are not randomly distributed but possess quantifiable statistical order, proposing $\overline{n_{+}}$ as a robust physical quantity to analyze the topological structure and complexity of chimera states.
@@
Introduction. Large systems of coupled oscillators are a fundamental model for understanding cooperative behaviors in a wide range of physical, chemical, and biological systems [Kuramoto1984, strogatz2003sync]. When the coupling is sufficiently weak, the complex dynamics can be simplified and described by a phase variable through phase reduction. Depending on the form of the coupling function, these oscillators exhibit not only global synchronization and desynchronization [RevModPhys.77.137], but also chimera states [kuramoto2002coexistence, PhysRevLett.93.174102], which are characterized by the spatial coexistence of coherent and incoherent regions. Initially found in non-locally coupled one-dimensional systems, chimera states have since been observed in two- and three-dimensional systems [PhysRevE.69.036213, PhysRevE.70.065201, article23, article33], as well as in populations of oscillators [PhysRevLett.101.084103]. They have been experimentally realized in chemical [PhysRevLett.110.244102, Totz2018, Tinsley2012], mechanical [doi:10.1073/pnas.1302880110, Kapitaniak2014], and electronic oscillators [PhysRevE.90.030902, PhysRevE.90.032905] and are particularly relevant to the study of neuronal networks [MAJHI2019100].
To investigate these phenomena, we employ the non-local Sakaguchi-Kuramoto model [10.1143/PTP.76.576] on a two-dimensional $N\times N$ square lattice, where each lattice point hosts an identical oscillator. Each oscillator interacts with $N(R)$ neighboring oscillators within a finite distance $R$. The governing equation is given by:
@@
In two dimensions, the variety of patterns is abundant, with spiral waves being the most archetypal example [PhysRevE.70.065201, PhysRevLett.104.044101]. The Kuramoto-like model on a two-dimensional lattice shares a powerful analogy with the XY model of planar spins. The study of two-dimensional Kuramoto-like models from a vortex perspective has recently gained traction, with several studies investigating vortex dynamics. These works have explored the Kuramoto-like model for specific cases, such as nearest-neighbor coupling ($R=1$) with random intrinsic frequencies and the entrainment transition [PhysRevE.82.036202], or through a duality transformation to expose underlying vortex structures [PhysRevE.103.032204] and in 2D Spin Torque Oscillators [Flovik_2016]. Additionally, research on the 2D-XY model of self-driven rotors [PhysRevLett.127.088004] has provided insights into defect dynamics and their superdiffusive behavior. Furthermore, the dynamics of traveling spiral wave chimeras and their phase-lag dependent transitions have also been studied in detail [bataille2023traveling]. However, to the best of our knowledge, the specific case of a non-local coupling ($R>1$) with a non-zero phase lag has not been studied from a topological perspective. This particular combination of parameters is crucial for the emergence of chimera states. The primary objective of this paper is to investigate these chimera states through a topological lens.
@@
Analytical Results. Following Ref. [PhysRevE.69.036213], we rewrite the governing equation (2) in the form of self-consistent equation
@@
$G(|\bm{x}-\bm{x}^{\prime}|)=\begin{cases}\frac{1}{\pi R^{2}},&\text{if }|\bm{x}-\bm{x}^{\prime}|<R,\\ 0,&\text{if }|\bm{x}-\bm{x}^{\prime}|>R.\end{cases}$ (5)
Following the approach in Ref. [PhysRevLett.104.044101], we adopt the Spiral Wave Ansatz [15cd7884-e7e8-3592-aee4-ba066c0c3e21]:
@@
The self-consistent equations remain analytically intractable. To proceed, a series expansion in the small parameter $\alpha$ is needed, following the method in Ref. [PhysRevLett.104.044101]:
$\Delta$ $=\Delta_{1}\alpha+O(\alpha^{2}),$ (9)
$A(r)$ $=A_{0}(r)+A_{1}(r)\alpha+O(\alpha^{2}),$
$\Psi(r)$ $=\Psi_{0}(r)+\Psi_{1}(r)\alpha+O(\alpha^{2}).$
@@
As $r\to 0$, the integral domain shrinks, $\lim_{r\to 0}A_{0}(r)=0.$. At $r=R$, the integration range becomes $0\leq u\leq 4R^{2}$. Substituting into the integrand yields $A_{0}(R)=\frac{8}{3\pi}$. In the limit $r\gg R$, $\lim_{r\to\infty}A_{0}(r)=1$ (see Supplementary Information S2 for details).
@@
The incoherent oscillators are defined by the condition $A(r)<|\Delta|$. Let $\rho$ be the radius of the incoherent core. We approximate $A_{0}(\rho)$ linearly as $A_{0}(\rho)\approx A_{0}^{\prime}(0)\rho=\Delta_{1}\alpha$. Since $A_{0}(R)=\frac{8}{3\pi}$ and the function is approximately linear for small $r$, we estimate the derivative as $A_{0}^{\prime}(0)\approx\frac{8}{3\pi R}$. Substituting this into the previous expression and using $\Delta_{1}\approx 1$, we find
$\rho\approx\frac{1}{A_{0}^{\prime}(0)}\alpha=\frac{3\pi}{8}R\alpha.$ (13)
@@
We also conduct a linear stability analysis for plane waves $\psi(\bm{r},t)=\Omega t+\bm{Q}\cdot\bm{r}$, where $\bm{Q}$ is the wave vector, $\Omega$ is the corresponding frequency (see Supplementary Information S1 for details). We find when $0<\alpha<\tfrac{1}{2}\pi$(the in-phase regime), the stable domain is determined to be in the approximate range of $0<\rho<2.2917$, where $\bm{\rho}=R\bm{Q}$.
@@
A single vortex: First, we investigate a single isolated vortex to validate our numerical results against the analytical solution in the small $\alpha$ limit. We employ $N=151,\ t=100,\ \alpha=10^{\circ}$ and impose reflection boundary conditions. A $\mathit{+1}$ defect is initialized at the center of the simulation box, as shown in Fig. 1(a) and (b). The profiles of the amplitude $|z|$ and phase $\Phi$ along the diagonal cross-section exhibit excellent agreement with the analytical predictions shown in Fig. 1(c) and (d). In the coherent core region, the phase $\Phi$ remains nearly constant before dropping abruptly at $r=R$, resulting in a discontinuous derivative. From Eq. (13), we obtain $\rho\approx 1.23$, which is smaller than distance $\rho_{1}=\sqrt{5/2}$ from the vortex center to the next-nearest grid point. This implies that the system can only form an incoherent core with just four grid points, which is in good agreement with the result in Fig. 1(a).
@@
We systematically investigate the spatial patterns that emerge in the system over the range of the phase lag $\alpha\in[0,\pi/2)$. Fig. 2 presents the phase distribution ($\phi$), the order parameter magnitude ($|z|$), and the diagonal profile of the complex order parameter at four representative values: $\alpha=2^{\circ},30^{\circ},45^{\circ}$, and $65^{\circ}$. We identify three primary dynamical regimes based on the patterns in the final state: (a),(b) a planar oscillation state, (c) a stable spiral wave state with incoherent cores, and (d) a turbulence-like state driven by the expansion of incoherent regions. To demarcate the parameter intervals for these distinct regimes, we define three critical values: $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$. These findings are in good agreement with previous studies, such as Ref. [PhysRevE.70.065201], which reported similar patterns arising from random initial conditions.
@@
In the regime $\alpha_{1}<\alpha<\alpha_{2}$ (with $\alpha_{2}\approx 63^{\circ}$), the system sustains a stable pair of spiral waves with persistent incoherent cores, as shown in Fig. 2(c). The order parameter distribution reveals an interference pattern between the spiral cores that forms a nearly square geometry, an artifact that we attribute to the square-shaped boundary conditions of the simulation, As $\alpha\approx 60^{\circ}$, the symmetry between the two spiral waves is progressively broken, and the system ultimately evolves into a state consisting of one large spiral wave and one (or more) smaller companion spiral.
@@
Defect Statistics. We conducted a statistical analysis of topological defects to elucidate their behavior under varying phase lag. As shown in Fig. 4, we plot the distribution of the total positive winding number $n_{+}$ for the phase lag $\alpha=45^{\circ}$ and $\alpha=55^{\circ}$. We first compare these data to the theoretical zero-truncated Poisson distribution given by:
$P(n_{+})=\frac{(\overline{n_{+}})^{n_{+}}e^{-\overline{n_{+}}}}{n_{+}!(1-e^{-\overline{n_{+}}})},\quad n_{+}\geq 1$ (14)
(see Supplementary Information S4 for details).
Figure 4: Statistics of the total positive winding number $n_{+}$ under different phase lags. (a) $\alpha=45^{\circ}$. (b) $\alpha=55^{\circ}$. The left panels show the histogram of $n_{+}$ compared to a truncated Poisson distribution [Eq. 14] and a binomial distribution [Eq. 15]. The right panels show the corresponding time evolution of $n_{+}$ (blue line) and its running time-average (red dashed line), $\langle n_{+}\rangle(\tau)=\frac{1}{\tau}\sum_{t=100}^{\tau}n_{+}(t)$ (Note that the $\alpha=55^{\circ}$ case is not compared with the inomial distribution, because $\sigma^{2}>\overline{n_{+}}$).
This model, however, inherently assumes that the variance of the distribution is equal to its mean ($\sigma^{2}=\overline{n_{+}}$). We find that our measured data exhibits a variance significantly smaller than its mean ($\sigma^{2}<\overline{n_{+}}$), which contradicts the Poisson assumption. This variance suppression ($\sigma^{2}<\overline{n_{+}}$) is the defining characteristic of the binomial distribution, suggesting a finite number of underlying ”sites” ($N_{\text{core}}$) capable of generating defects. Therefore, we also compare the data to a binomial distribution, $B(N_{\text{core}},p)$, whose probability mass function is:
$P(n_{+})=\binom{N_{\text{core}}}{n_{+}}p^{n_{+}}(1-p)^{N_{\text{core}}-n_{+}}.$ (15)
(see Supplementary Information S4 for details). Crucially, the parameters $N_{\text{core}}$ and $p$ are not arbitrary fitting parameters. They are determined directly from the measured mean $\overline{n_{+}}$ and variance $\sigma^{2}$ of the data segment via the relations:
$N_{\text{core}}=\frac{(\overline{n_{+}})^{2}}{\overline{n_{+}}-\sigma^{2}}\quad\text{and}\quad p=\frac{\overline{n_{+}}}{N_{\text{core}}}.$ (16)
@@
To quantitatively assess the influence of $\alpha$ on the system properties, we computed the corresponding $\overline{n_{+}},\sigma$, and the information entropy $S_{\mathrm{Shannon}}=-\sum_{n_{+}=1}^{\infty}P(n_{+})\ln P(n_{+})$ determined by Eq. (14). We first performed a linear fit for $\ln\overline{n_{+}}$ and $\ln\sigma$ versus $\alpha$, as shown in Fig. 5(a). The results show strong linear correlations, with coefficients of determination $R^{2}$ exceeding 0.99. The fitting results are as follows: $\ln\overline{n_{+}}=12.75\alpha/\pi-0.43,\ln\sigma^{2}=14.96\alpha/\pi-1.51,$ which implies an exponential form:
$\overline{n_{+}}=ae^{b\alpha},\quad\sigma^{2}=ce^{d\alpha},$ (17)
@@
$N_{\text{core}}(\alpha)=\frac{a^{2}e^{2b\alpha}}{ae^{b\alpha}-ce^{d\alpha}}$ (18)
This formula predicts super-exponential growth and divergence of the core size for increasing $\alpha$ (note that $b<d$). In fact, $\sigma^{2}$ is deviated from the exponential growth law and exceeds $\overline{n}_{+}$ at $\alpha=55^{\circ}$, where $N_{\text{core}}$ diverges according to Eq.(16) and the binomial statistics break down.
To further verify the exponential relationship, we repeated the above analysis for different coupling radii $R=4,8$. We can observe that the influence of $R$ on the slope of the fitted curve is comparatively small. The intercept ($\overline{n}_{+}$ at $\alpha=0$) increases with $R$, but the dependence is milder than the scaling of the area of the coupling range ($\propto R^{2}$); for example, the $\overline{n_{+}}$ for $R=4$ and $R=8$ are different by the ratio $3.4$, which is smaller than $(8/4)^{2}=4$.
We also examined the effect of system size by repeating the analysis at $N=150$ and the fitting curve remains nearly unchanged, indicating that the system size has little effect on the conclusion. The relationship appears to be universal.
Figure 5: (a) Dependence of the average total positive winding number $\overline{n_{+}}$ (left axis) and its standard deviation $\sigma$ (right axis) on the phase lag $\alpha$. The red markers show $\ln(\overline{n_{+}})$ for different interaction radii. The blue markers show $\ln(\sigma)$ for $R=6$. The lines represent linear fits to the corresponding data points. (b) Comparison of the measured entropy with the theoretical binomial and Poisson entropy.
@@
It is observed that for $\alpha<55^{\circ}$, the measured entropy agrees remarkably well with the binomial prediction. Around $\alpha\approx 50^{\circ}$, the system exhibits a sharp rise in entropy, reaching and even surpassing the level expected from a Poisson process. We identify this point as a new critical value, $\alpha^{*}$, marking a transition from binomial‐like to Poisson‐like behavior. This transition indicates that the system enters a highly unstable and fluctuation-dominated regime.
Discussion. This study reveals the statistical patterns of topological defects in spiral wave chimera states within two-dimensional phase oscillator networks, as well as the dependence of the total positive winding number on the phase lag. In stable chimera states, we observe that the average total positive winding number $\overline{n_{+}}$ exhibits a clear exponential growth relationship with the phase lag $\alpha$, following $\overline{n_{+}}=ae^{b\alpha}$. This relationship is concise in form and fits the data well. The phenomenon demonstrates robustness within a specific parameter range and is reproduced under various conditions, suggesting that topological excitations exhibit an underlying statistical order. Furthermore, the shape of histogram is closely correlated with the stability of the spiral waves. This confirms that the topological structure in chimera states is not entirely disordered, but rather reflects the system’s intrinsic properties in an organized manner. These results indicate that topological defects in spiral wave chimera states are not randomly distributed. Instead, they contain observable and quantifiable structural features that evolve regularly with system parameters. Therefore, the average positive winding number may serve as a new physical quantity for quantitatively analyzing the topological structure and complexity of chimera states.
Our study suggests the existence of different scaling relationships depending on the value of $\alpha$. The perturbation analysis in Section II demonstrated that in the limit $\alpha\to 0$, the incoherent core radius $\rho$ grows linearly with $\alpha$ ($\rho\propto\alpha$). This result describes the stable, geometric growth of the core in the small-$\alpha$ regime. In contrast, the central finding of this paper—the exponential growth of the average positive winding number ($\overline{n_{+}}\propto e^{b\alpha}$)—is a statistical law observed in the stable chimera regime where $\alpha$ is relatively large ($\alpha_{1}<\alpha<\alpha_{2}$). This transition from linear to exponential scaling may signify a crossover or bifurcation, separating two distinct physical mechanisms. At small $\alpha$, the system is dominated by the geometric expansion of the core, whereas at larger $\alpha$ (in the stable chimera regime), the dominant mechanism becomes the active topological excitation—the continuous generation and annihilation of defect pairs—within that core. At $\alpha=\alpha^{*}$, marking a transition from binomial‐like to Poisson‐like behavior. We interpret this statistical transition physically. The binomial-like distribution suggests a constrained regime, where defect generation is limited. The Poisson-like distribution, conversely, signifies an unconstrained regime, where defects behave as independent, random excitations (the continuous generation and annihilation of pairs). This shift from a constrained to unconstrained state strongly parallels the binding-unbinding BKT transition observed in the 2D XY model. A detailed investigation of how one scaling regime transitions to the other, near the critical values $\alpha_{1}$ and $\alpha_{2}$, remains a crucial future task. Such a study could provide key insights into the mechanisms that stabilize and destabilize the spiral wave chimera state.
Acknowledgements.
The author acknowledges financial support from the AGS RISE Program (Recruitment Strategic Proposal Type Program) at Tohoku University.

### v2

@@ Abstract
Chimera states, characterized by the coexistence of coherent and incoherent domains, represent a paradigm of self-organization in complex systems. In this study, we introduce a topological analysis method based on winding numbers to characterize the dynamics of spiral wave chimeras in a two-dimensional phase oscillator network. Our investigation reveals distinct scaling laws governing the system’s evolution across the phase lag $\alpha$. Perturbation analysis in the limit $\alpha\to 0$ demonstrates that the incoherent core radius scales linearly with $\alpha$. In contrast, within the stable chimera regime, the average total positive winding number $\mu$ follows a clear exponential growth law $\mu=ae^{b\alpha}$. This scaling disparity signals a physical crossover from a regime dominated by geometric core expansion to one driven by active topological excitation. Furthermore, we identify a statistical transition in the defect distribution from binomial-like to Poisson-like behavior at a critical threshold $\alpha^{*}$. These results demonstrate that topological defects possess intrinsic statistical order, establishing $\mu$ as a robust macro-variable for analyzing the structural complexity of chimera states.
@@
Introduction. Large systems of coupled oscillators are a fundamental model for understanding cooperative behaviors in a wide range of physical, chemical, and biological systems [11, 26]. When the coupling is sufficiently weak, the complex dynamics can be simplified and described by a phase variable through phase reduction. Depending on the form of the coupling function, these oscillators exhibit not only global synchronization and desynchronization [3], but also chimera states [10, 2], which are characterized by the spatial coexistence of coherent and incoherent regions. Initially found in non-locally coupled one-dimensional systems, chimera states have since been observed in two- and three-dimensional systems [25, 9, 20, 15, 12], as well as in populations of oscillators [1]. They have been experimentally realized in chemical [19, 28, 27], mechanical [18, 8], and electronic oscillators [21, 7] and are particularly relevant to the study of neuronal networks [16].
To investigate these phenomena, we employ the non-local Sakaguchi-Kuramoto model [23] on a two-dimensional $N\times N$ square lattice, where each lattice point hosts an identical oscillator. Each oscillator interacts with $N(R)$ neighboring oscillators within a finite distance $R$. The governing equation is given by:
@@
In two dimensions, a rich variety of patterns can emerge, among which spiral wave chimeras are the most archetypal example [9, 17] and have been observed experimentally [29]. The Kuramoto-like model on a two-dimensional lattice shares a powerful analogy with the XY model of planar spins. The study of two-dimensional Kuramoto-like models from a vortex perspective has recently gained traction, with several studies investigating vortex dynamics. These works have explored the Kuramoto-like model for specific cases, such as nearest-neighbor coupling ($R=1$) with random intrinsic frequencies and the entrainment transition [13], or through a duality transformation to expose underlying vortex structures [24] and in 2D Spin Torque Oscillators [6]. Additionally, research on the 2D-XY model of self-driven rotors [22] has provided insights into defect dynamics and their superdiffusive behavior. Furthermore, the dynamics of traveling spiral wave chimeras and their phase-lag dependent transitions have also been studied in detail [4]. However, to the best of our knowledge, the specific case of a non-local coupling ($R>1$) with a non-zero phase lag has not been studied from a topological perspective. This particular combination of parameters is crucial for the emergence of chimera states. The primary objective of this paper is to investigate these chimera states through a topological lens.
@@
Analytical Results. Following Ref. [25], we rewrite the governing equation (2) in the form of self-consistent equation
@@
$G(|\bm{x}-\bm{x}^{\prime}|)=\cases{}{1}{\pi R^{2}},&\text{if }|\bm{x}-\bm{x}^{\prime}|<R,\\ 0,&\text{if }|\bm{x}-\bm{x}^{\prime}|>R.$ (5)
Following the approach in Ref. [17], we adopt the Spiral Wave Ansatz [5]:
@@
The self-consistent equations remain analytically intractable. To proceed, a series expansion in the small parameter $\alpha$ is needed, following the method in Ref. [17]:
$aligned\Delta&=\Delta_{1}\alpha+O(\alpha^{2}),\\ A(r)&=A_{0}(r)+A_{1}(r)\alpha+O(\alpha^{2}),\\ \Psi(r)&=\Psi_{0}(r)+\Psi_{1}(r)\alpha+O(\alpha^{2}).$ (9)
@@
As $r\to 0$, the integral domain shrinks, $\lim_{r\to 0}A_{0}(r)=0.$. At $r=R$, the integration range becomes $0\leq u\leq 4R^{2}$. Substituting into the integrand yields $A_{0}(R)=\frac{8}{3\pi}$. In the limit $r\gg R$, $\lim_{r\to\infty}A_{0}(r)=1$ (see Supplementary Information S1 for details).
@@
The incoherent oscillators are defined by the condition $A(r)<|\Delta|$. Let $r_{\text{core}}$ be the radius of the incoherent core. We approximate $A_{0}(r_{\text{core}})$ linearly as $A_{0}(r_{\text{core}})\approx A_{0}^{\prime}(0)r_{\text{core}}=\Delta_{1}\alpha$. Since $A_{0}(R)=\frac{8}{3\pi}$ and the function is approximately linear for small $r$, we estimate the derivative as $A_{0}^{\prime}(0)\approx\frac{8}{3\pi R}$. Substituting this into the previous expression and using $\Delta_{1}\approx 1$, we find
$r_{\text{core}}\approx\frac{1}{A_{0}^{\prime}(0)}\alpha=\frac{3\pi}{8}R\alpha.$ (13)
@@
To evaluate the robustness of these patterns, we perform a linear stability analysis of the plane wave solutions [14, 9], defined by $\psi(\bm{r},t)=\Omega t+\bm{Q}\cdot\bm{r}$, where $\bm{Q}$ denotes the wave vector and $\Omega$ represents the corresponding frequency (see Supplementary Information S2 for the full derivation). The local wave vector of the spiral arms is estimated via the phase gradient:
$\bm{Q}=\nabla\Phi=\frac{d\Psi(r)}{dr}\hat{\bm{r}}+\frac{1}{r}\hat{\bm{\theta}}$
Our analysis clarifies how the coupling range $R$ and the phase lag $\alpha$ synergistically dictate the stability of the spiral wave arms. In the in-phase regime ($0<\alpha<\frac{\pi}{2}$), the stability domain is constrained to the approximate range $0<\rho<2.2917$, where $\rho=R|\bm{Q}|$. Conversely, in the anti-phase regime ($\frac{\pi}{2}<\alpha<\pi$), the stable domain shifts to $4.2005<\rho<6.1876$. Notably, as the coupling range $R$ increases, the stability criterion is more readily violated, which corresponds to a progressive shortening of the spiral arm wavelength. Consequently, within the regime $\frac{\pi}{2}<\alpha<\pi$, the spiral wave configurations often become intrinsically unstable, leading to the emergence of more complex spatiotemporal patterns.
@@
A single vortex: First, we investigate a single isolated vortex to validate our numerical results against the analytical solution in the small $\alpha$ limit. We employ $N=151,\ t=100,\ \alpha=10^{\circ}$ and impose reflective boundary conditions. A $\mathit{+1}$ defect is initialized at the center of the simulation box, as shown in Fig. 1(a) and (b). The profiles of the amplitude $|z|$ and phase $\Phi$ along the diagonal cross-section exhibit excellent agreement with the analytical predictions shown in Fig. 1(c) and (d). In the coherent core region, the phase $\Phi$ remains nearly constant before dropping abruptly at $r=R$, resulting in a discontinuous derivative. From Eq. (13), we obtain $\rho\approx 1.23$, which is smaller than distance $\rho_{1}=\sqrt{5/2}$ from the vortex center to the next-nearest grid point. This implies that the system can only form an incoherent core with just four grid points, which is in good agreement with the result in Fig. 1(a).
@@
We systematically investigate the spatial patterns that emerge in the system over the range of the phase lag $\alpha\in[0,\pi/2)$. Fig. 2 presents the phase distribution ($\phi$), the order parameter magnitude ($|z|$), and the diagonal profile of the complex order parameter at four representative values: $\alpha=2^{\circ},25^{\circ},45^{\circ}$, and $65^{\circ}$. We identify three primary dynamical regimes based on the patterns in the final state: (a),(b) a planar oscillation state, (c) a stable spiral wave state with incoherent cores, and (d) a turbulence-like state driven by the expansion of incoherent regions. To demarcate the parameter intervals for these distinct regimes, we define three critical values: $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$. These findings are in good agreement with previous studies, such as Ref. [9], which reported similar patterns arising from random initial conditions.
@@
In the regime $\alpha_{1}<\alpha<\alpha_{2}$ (with $\alpha_{2}\approx 63^{\circ}$), the system sustains a stable pair of spiral waves with persistent incoherent cores, as shown in Fig. 2(c). The order parameter distribution reveals an interference pattern between the spiral cores that forms a nearly square geometry, an artifact that we attribute to the square-shaped boundary conditions of the simulation. As $\alpha\approx 60^{\circ}$, the symmetry between the two spiral waves is progressively broken, leading the system to settle into a configuration characterized by a dominant large spiral wave accompanied by one or more smaller spirals. Furthermore, during the translation of these spiral wave chimeras, distinct filamentary structures emerge within the incoherent core [4].
@@
Defect Statistics. We conducted a statistical analysis of topological defects to elucidate their behavior under varying phase lag (see Supplementary Information S4 for details). As shown in Fig. 4, we plot the distribution of the total positive winding number $n_{+}$ for the phase lag $\alpha=45^{\circ}$ and $\alpha=55^{\circ}$. We first compare these data to the theoretical zero-truncated Poisson distribution given by:
$P(n_{+})=\frac{{\mu}^{n_{+}}e^{-{\mu}}}{n_{+}!(1-e^{-{\mu}})},\quad n_{+}\geq 1$ (14)
where ${\mu}\equiv\overline{n_{+}}$ .
Figure 4: Statistics of the total positive winding number $n_{+}$ under different phase lags. (a) $\alpha=45^{\circ}$. (b) $\alpha=55^{\circ}$. The left panels show the histogram of $n_{+}$ compared to a truncated Poisson distribution [Eq. 14] and a binomial distribution [Eq. 15]. The right panels show the corresponding time evolution of $n_{+}$ (blue line) and its running time-average (red dashed line), $\langle n_{+}\rangle(\tau)=\frac{1}{\tau}\sum_{t=100}^{\tau}n_{+}(t)$ (Note that the $\alpha=55^{\circ}$ case is not compared with the binomial distribution, because $\sigma^{2}>\mu$).
This model, however, inherently assumes that the variance of the distribution is equal to its mean ($\sigma^{2}={\mu}$). We find that our measured data exhibits a variance significantly smaller than its mean ($\sigma^{2}<{\mu}$), which contradicts the Poisson assumption. This variance suppression ($\sigma^{2}<{\mu}$) is the defining characteristic of the binomial distribution, suggesting a finite number of underlying sites ($N_{\text{core}}$) capable of generating defects. Therefore, we also compare the data to a binomial distribution, $B(N_{\text{core}},p)$, whose probability mass function is:
[…]
Crucially, the parameters $N_{\text{core}}$ and $p$ are not arbitrary fitting parameters. They are determined directly from the measured mean $\mu$ and variance $\sigma^{2}$ of the data segment via the relations:
$N_{\text{core}}=\frac{\mu^{2}}{\mu-\sigma^{2}}\quad\text{and}\quad p=\frac{\mu}{N_{\text{core}}}.$ (16)
@@
To quantitatively assess the influence of $\alpha$ on the system properties, we computed the corresponding $\mu,\sigma$, and the information entropy $S_{\mathrm{Shannon}}=-\sum_{n_{+}=1}^{\infty}P(n_{+})\ln P(n_{+})$ determined by Eq. (14). We first performed a linear fit for $\ln\mu$ and $\ln\sigma$ versus $\alpha$, as shown in Fig. 5(a). The results show strong linear correlations, with coefficients of determination exceeding 0.99. The fitting results are as follows: $\ln\mu=12.75\alpha/\pi-0.43,\ln\sigma^{2}=14.96\alpha/\pi-1.51,$ which implies an exponential form:
$\mu=ae^{b\alpha},\quad\sigma^{2}=ce^{d\alpha},$ (17)
@@
$N_{\text{core}}(\alpha)=\frac{a^{2}e^{2b\alpha}}{ae^{b\alpha}-ce^{d\alpha}}.$ (18)
This formula predicts super-exponential growth and divergence of the core size for increasing $\alpha$ (note that $b<d$). In fact, $\sigma^{2}$ is deviated from the exponential growth law and exceeds $\mu$ at $\alpha=55^{\circ}$, where $N_{\text{core}}$ diverges according to Eq.(16) and the binomial statistics break down. To verify the validity of the analytical expression for $N_{\text{core}}$, we performed numerical simulations to measure the core area. The comparison between the numerical data and the theoretical curve is illustrated in Fig. 5(b). we find that the theoretical node count and the simulated area follow the scaling relationship: $N_{\text{core}}\approx\frac{1}{k}(\pi r_{\text{core}}^{2})$ where the scaling factor $k\approx 3.02$. This close agreement across the surveyed range of $\alpha$ demonstrates that the evolution of the spiral wave core size is quantitatively governed by the exponential competition between the statistical parameters $\mu$ and $\sigma^{2}$.
To further verify the exponential relationship, we repeated the above analysis for different coupling radii $R=4,8$. We can observe that the influence of $R$ on the slope of the fitted curve is comparatively small. The intercept ($\mu$ at $\alpha=0$) increases with $R$, but the dependence is milder than the scaling of the area of the coupling range ($\propto R^{2}$); for example, the $\mu$ for $R=4$ and $R=8$ are different by the ratio $3.4$, which is smaller than $(8/4)^{2}=4$.
We also examined the effect of system size by repeating the analysis at $N=150,200$ and for the single-spiral case (see Supplementary Information S5 for details). The fitting curve remains nearly unchanged, indicating that the system size has little effect on the conclusion. The relationship appears to be universal.
Figure 5: (a) Dependence of the average total positive winding number $\mu$ and its standard deviation $\sigma^{2}$ on the phase lag $\alpha$. The red markers show $\mu$ for different interaction radii. The blue markers show $\ln(\sigma^{2})$ for $R=6$. The lines represent linear fits to the corresponding data points. (b) The numerical simulation results for the core area $\pi r_{\text{core}}^{2}$ are compared with the theoretical prediction $N_{\text{core}}(\alpha)$ [Eq. (18)]. (c) Comparison of the measured entropy with the theoretical binomial and Poisson entropy. (d) Comparison of spatial phase patterns near the critical transition point $\alpha^{*}$ at $t=190$, observed in the region $x,y\in[45,99]$ for $\alpha=54^{\circ}$ (top) and $\alpha=55^{\circ}$ (bottom).
@@
It is observed that for $\alpha<55^{\circ}$, the measured entropy agrees remarkably well with the binomial prediction. Around $\alpha\approx 55^{\circ}$, the system exhibits a sharp rise in entropy, reaching and even surpassing the level expected from a Poisson process. We identify this point as a new critical value, $\alpha^{*}$, marking a transition from binomial‐like to Poisson‐like behavior.
To elucidate the structural origin of this transition, we compare the spatial patterns at $\alpha=54^{\circ}$ and $\alpha=55^{\circ}$ in Fig. 5(d). Notably, at $\alpha=55^{\circ}$, the core exhibits the emergence of filamentous structures. These structural features and the identified transition point are in good agreement with previous studies on the transition from static to moving chimera states [4](see Supplementary Movie 1 for a dynamic visualization of this process). However, our study distinguishes itself in two key aspects: In previous works, the transition point is often distributed or broadened due to the inhomogeneity of natural frequencies, whereas our system of identical oscillators yields a sharper definition of the transition. While previous coarse-grained approaches could not resolve fine-scale dynamics, our method tracks individual topological defects. Consequently, our analysis provides a clearer and more quantitative description of the static-moving chimera transition.
Discussion. This study reveals the statistical laws governing topological defects in spiral wave chimera states and establishes the total positive winding number, $\mu$, as a robust macro-variable for characterizing these complex states. Our analysis leads to two major physical insights regarding the scaling, and the statistical transition of the chimera core.
First, we establish that in the stable chimera regime, $\mu$ follows a clear exponential growth with the phase lag ($\mu\sim e^{b\alpha}$). This relationship is concise in form and fits the data well. The phenomenon demonstrates robustness within a specific parameter range and is reproduced under various conditions, suggesting that topological excitations exhibit an underlying statistical order. Crucially, our scaling analysis refutes the hypothesis of simple geometric growth. The observed divergence between defect generation and core area expansion indicates that the defect population is driven by intrinsic dynamical instability rather than passive geometric scaling.
Second, we identified a critical transition point at $\alpha^{*}\approx 55^{\circ}$, marking a fundamental shift in the nature of the core turbulence. For $\alpha<\alpha^{*}$, the defect statistics follow a Binomial distribution, suggesting a constrained regime where defect generation is limited by steric repulsion or core capacity. However, as $\alpha$ exceeds $\alpha^{*}$, we observe a sharp rise in entropy and a crossover to a Poisson-like distribution, indicating an unconstrained regime dominated by independent, random excitations. Physically, this statistical transition coincides with the emergence of filamentous structures within the core and signals the onset of the moving chimera state.
In conclusion, topological defects in chimera states are not random artifacts but possess quantifiable structural features that evolve regularly with system parameters. The transition from the geometry-dominated linear growth at small $\alpha$ to the instability-driven exponential growth at larger $\alpha$, and finally to the Poissonian structural breakdown at $\alpha^{*}$, delineates the rich dynamical landscape of chimera states. A detailed investigation of how one scaling regime transitions to the other, near the critical values $\alpha_{1}$ and $\alpha_{2}$, remains a crucial future task. Such a study could provide key insights into the mechanisms that stabilize and destabilize the spiral wave chimera states.